OpenStudy (anonymous):

plzzzz need help ODEs \((\cos x) (e^y -y) \frac{dy}{dx}=e^y\sin 2x \) i did these steps but i couldnt continue separate variables \(\Large (1 - ye^{-y})dy = \frac{2\sin(x)\cos(x)}{\cos(x)}dx\) @ganeshie8 @terenzreignz

OpenStudy (usukidoll):

errr you could just cancel out the cosx and then take the antiderivative on both sides to see what you have

OpenStudy (anonymous):

Why not just integrate?

OpenStudy (anonymous):

\[ \int 2\sin(x)\;dx = -2\cos(x)+C \]

OpenStudy (usukidoll):

oh sorry yeah .. integrate... XD

OpenStudy (usukidoll):

yeah that's right integrate both sides... integration by parts for y?!

OpenStudy (ikram002p):

@BSwan check it out !

OpenStudy (anonymous):

by exact ??

OpenStudy (usukidoll):

exact? for real? that has to be in the form of M (x,y) dy + N (x,y) dx = 0

OpenStudy (usukidoll):

then you have to take My and Nx and if they are not the same, you need to find the integrating factor... still won't work? then exact isn't a good choice

OpenStudy (anonymous):

sin2x=2sinxcosx, so \((e^y-y)dy/dx=2\sin x e^y\) separating variables: \(1-ye^{-y}dy=2sinxdx\)

OpenStudy (ankitshaw):

\[(1-y/e^y)dy = 2sinxdx \] now u can use substitution method in LHS to solve it..